# Maxwell’s Demon at Work: Efficient Pruning by Leveraging Saturation of Neurons

### v1

@@ Abstract
When training deep neural networks, the phenomenon of dying neurons —units that become inactive or saturated, output zero during training— has traditionally been viewed as undesirable, linked with optimization challenges, and contributing to plasticity loss in continual learning scenarios. In this paper, we reassess this phenomenon, focusing on sparsity and pruning. By systematically exploring the impact of various hyperparameter configurations on dying neurons, we unveil their potential to facilitate simple yet effective structured pruning algorithms. We introduce Demon Pruning (DemP), a method that controls the proliferation of dead neurons, dynamically leading to network sparsity. Achieved through a combination of noise injection on active units and a one-cycled schedule regularization strategy, DemP stands out for its simplicity and broad applicability. Experiments on CIFAR10 and ImageNet datasets demonstrate that DemP surpasses existing structured pruning techniques, showcasing superior accuracy-sparsity tradeoffs and training speedups. These findings suggest a novel perspective on dying neurons as a valuable resource for efficient model compression and optimization.
Machine Learning, ICML
@@ 1 Introduction
Dying neurons, a phenomenon frequently observed during the learning process of neural networks, are traditionally viewed as detrimental, often leading to suboptimal performance (Maas et al., 2013; Xu et al., 2015) or loss of plasticity, especially in non-stationary settings (Lyle et al., 2023; Nikishin et al., 2022; Abbas et al., 2023). In response, multiple mitigation strategies have been proposed, from alternative activation functions without a hard-saturated state – such as Leaky ReLU (Maas et al., 2013), Swish (Ramachandran et al., 2018), GELU (Hendrycks & Gimpel, 2016) – to wide use of normalization (Lyle et al., 2023) or targeted resets of weights (D’Oro et al., 2022; Dohare et al., 2021).
In this work, we revisit this phenomenon through the lens of network sparsity and pruning. Drawing on both intuitive and theoretical insights, we showcase how hyperparameter choices can encourage the saturation of units, offering an effective strategy for pruning the architecture. We introduce Demon Pruning (DemP), which combines regularization and noise injection on active units to control the proliferation of dead neurons. Drawing an analogy to Maxwell’s demon thought experiment  (Maxwell, 1872) in thermodynamics, DemP leverages the asymmetry observed in the saturation process, guiding units from an active to an inactive state.
The originality of DemP lies in its active promotion of unit saturation, dynamically yielding sparser subnetworks throughout training. Its simplicity, allowing dead units to be directly identified by inspecting their activations with a few forward passes, ensures pruning at nearly no additional cost, in contrast with many existing methods (see e.g., Lee et al., 2023). Despite its simplicity, DemP achieves superior accuracy-sparsity tradeoffs when compared to strong comparable baselines like EarlyCrop (Rachwan et al., 2022) or SNAP (Verdenius et al., 2020). For instance, it outperforms by up to $\sim 2.5\%$ beyond $80\%$ sparsity training a ResNet-18 on CIFAR-10 or a ResNet-50 on ImageNet, coupled with a higher training speedup, reaching up to 1.23x faster on ImageNet. Finally, beyond demonstrating strong empirical performance, DemP seamlessly integrates into any training algorithm and can be readily combined with existing pruning techniques.
Our primary contributions are as follows:
@@
    Insights into Neuron Saturation. We delve into the mechanisms of saturating units, shedding light on the significant role played by stochasticity and the impact of key hyperparameters like learning rate, batch size, and regularization parameters (Sec. 2).
@@
    A Structured Pruning Method. Building on our insights, we present DemP, a dynamic sparsity approach that encourages a controlled proliferation of dead neurons and removes them in real-time as they arise during training, resulting in substantial training speedups (Sec. 3)
@@
    Empirical Validation. Extensive experiments on CIFAR-10 and ImageNet datasets demonstrate that DemP, despite its simplicity and versatility, surpasses strong existing structured pruning baselines in terms of accuracy-compression tradeoffs and achieves results comparable to unstructured pruning methods (Sec. 4).
@@ 2 Saturating Units: An Analysis
The goal of this section is to offer theoretical insights into the occurrence of dead neurons and explore the impact of different training heuristics and hyperparameters.
Given a neural network and a set of $n$ training data samples, we denote by $a^{\ell}_{j}\in\mathbb{R}^{n}$ the vector of activations of the $j$th neuron in layer $\ell$ for each training input. We adopt the following definition of a dead neuron throughout the paper:
[…]
Many modern architectures use activations functions with a saturation region that typically includes 00 at its boundary. In this case, when a neuron becomes inactive during training, its incoming weights also receive zero—or very small333e.g., due to regularization.— gradients, which makes it difficult for the neuron to recover. In this paper, we mostly work with the Rectified Linear Unit (ReLU) activation function, $\sigma(x)=\max(0,x)$. In this case, the activity of a neuron depends on the sign of the corresponding pre-activation feature.
@@ 2.1 Neurons Die During Training
We begin with some empirical observations. Applying the above criterion (Footnote 1) with threshold parameter $\epsilon=0.01$, we monitor the accumulation of dead neurons during training of a Resnet-18 (He et al., 2016) on CIFAR-10 (Krizhevsky et al., 2009) with the Adam optimizer (Kingma & Ba, 2015), with various learning rates and choices of activation functions.
@@
Intuition.
Similar to Maxwell’s demon thought experiment (Maxwell, 1872), one can picture an asymmetric behavior of how weights can move across the boundary that delimitates the saturated versus non-saturated state of a unit, akin to the demon’s selective passage of molecules. Neurons, or more specifically their weights, can move freely within the active region, but once they enter the inactive region their movement is impeded, similar to how the demon can trap molecules in a low-energy subspace. In this context, neuron death can be influenced by various factors, e.g., noise from the data, being too close to the border, or taking too large gradient steps. Once the neuron moves into the inactive zone, neurons can only be reactivated if the boundary itself shifts. This asymmetry makes it more likely for neurons to die than to revive.
We formalize this analogy with simple theoretical models in Appendix B. These models are meant to capture the multiplicative (i.e. parameter dependent) nature of the gradient noise in SGD. Multiplicative noise in stochastic processes is known to cause regions with lower noise magnitude to act as attractors (Oksendal, 2010). Essentially, lower noise regions (such as the inactive region of a saturating unit) tend to retain the system due to reduced impact, while higher noise regions push it away. We explicitly demonstrate this phenomenon in the theoretical settings of Appendix B.
Figure 1: Dead neuron accumulation for a ResNet-18 trained on CIFAR-10 with different activation functions and values of the learning rate. We use a negative slope of $\alpha=0.05$ for Leaky ReLU and a $\beta=1$ for Swish.
@@ 2.2 Factors Impacting Dying Ratios
The above intuition suggests that maintaining activations close to zero increases the likelihood of neuron death, aligning them closer to the inactive region (the negative domain for ReLU networks). As illustrated in Fig. 2, training a ResNet-18 on CIFAR 10 (with Adam or SGD+momentum) with a higher L2 or Lasso regularization leads to sparser networks caused by an accumulation of dead units. Section 3 elaborates on the approach with DemP, where we opt to apply regularization exclusively on the scale parameters of the normalization layers, following Liu et al. (2017a). This tends to better preserve performance and yields significantly improved accuracy-sparsity tradeoffs, as shown in Fig 6 in Appendix F.
Role of noise.
Our simple model suggests a pivotal role played by the noise of the training process in the occurrence of dying neurons. To investigate this in a simple setting, we train small MLPs on a subset of 10 000 images of the MNIST dataset in three different noisy regimes: vanilla SGD, pure SGD noise (obtained by isolating the noisy part of the minibatch gradient), and pure Gaussian noise. As observed in Fig. 10 in Appendix D.1, while pure SGD noise training yields dying ratios comparable to SGD, training with pure Gaussian noise is much less prone to dead neuron accumulation. We hypothesize that this difference is due to the asymmetry of SGD noise: since the gradients are 0 for dead neurons, only (the weights of) live neurons are subject to noisy updates. In contrast, the weights of inactive neurons get updated under Gaussian noise training, which increases their probability of recovery. As we spell out in Section 3, DemP injects Gaussian noise exclusively to weights of live neurons during training, which reveals effective to encourage neuron death (see Fig. 3 and Fig. 10, middle plot).
Figure 2: Increased regularization during training increases the ratio of dead units, as showcased here for a ResNet-18 trained on CIFAR-10. We use $\cdot(\gamma)$ to denote when regularization is applied solely to the scale parameters of the normalization layers.
@@
Figure 3: Augmenting training updates with asymmetric Gaussian noise, sampled from $\mathcal{N}(x|0,\sigma^{2})$ and applied to live neurons only, also leads to higher levels of dead unit accumulation, as showcased here for a ResNet-18 trained on CIFAR-10 with various values of the Lasso($\gamma$) regularization parameter.
@@
We note that other ways to control the SGD noise variance include varying the learning rate and/or the batch size. Fig. 11 (right) in Appendix D.1 illustrates how increasing the learning rate or diminishing the batch size indeed increases dying ratios. However, it proved more challenging to balance accuracy and sparsity by adjusting these hyperparameters, as they also significantly affect overall performance. Instead, DemP maintains these parameters at their default or task-optimized values, facilitating its integration into existing training routines.
Optimizer.
@@
We expect the choice of optimizer to affect the final count of dead neurons post-training (e.g. for adaptive optimizers, by altering the effective learning rate per parameter). Notably, we observe a significant discrepancy between using the Adam optimizer (Kingma & Ba, 2015) and SGD with momentum (SGDM) (Fig. 2). As also emphasized by Lyle et al. (2023), we attribute this difference primarily to the specific hyperparameter selection for Adam ($\beta_{1},\beta_{2},\epsilon$), which has a substantial impact on neuron death (further discussed in Appendix E). Results from experiments with both optimizers are presented in the next section.
@@
Figure 4: Top: For ResNet-18 networks on CIFAR-10 trained with Adam (ReLU), DemP can find sparser solutions maintaining better performance than other structured approaches. Higher sparsities with DemP are obtained by increasing the peak strength of the added scheduled regularization. Bottom: With SGDM, DemP performance is on par, without significant differences between methods. Left: Neural sparsity, structured methods. Right: Weight sparsity, structured methods.

### v2

@@ Abstract
[…]
@@ 1 Introduction
In neural network training, dying neurons—those that saturate or become inactive during the learning process—have long been considered undesirable (Maas13rectifiernonlinearities; xu2015empirical), leading to performance degradation or loss of plasticity, especially in non-stationary settings (DBLP:conf/icml/LyleZNPPD23; nikishin2022primacy; DBLP:journals/corr/AbbasZMWM2023_PlasticityLossInDRL). Past research has proposed solutions such as alternative activation functions like Leaky ReLU or Swish (Maas13rectifiernonlinearities; DBLP:conf/iclr/RamachandranZL18), or methods to reset weights (d2022sample; arXiv2021_DohareRS). However, the role of dead neurons remains poorly understood.
In this work, we reframe the phenomenon of neuron death and saturation as a resource for pruning, allowing us to reduce the size of neural networks dynamically during training. Inspired by Maxwell’s Demon, which selectively filters molecules in thermodynamics  (maxwell1872theory), we introduce Demon Pruning (DemP), a sparse training method that (i) promotes neuron death by applying asymmetric Gaussian noise to the weights of active neurons and using a one-cycle regularization schedule, and (ii) iteratively removes dead neurons during training. This approach enables highly structured sparsity with minimal performance loss.
While DemP builds upon simple regularization-based methods (LessIsMoreZhouAP16; PruneTrainLymCZWSE19; liu2017learning), unlike prior approaches passively inducing sparsity, DemP actively promotes unit saturation during training, yielding progressively sparser subnetworks. Compared to existing methods (jaxpruner), a key feature of DemP is the simplification of the pruning process: by inspecting neuron activations with just a few forward passes, dead units are easily identified and pruned at minimum computational costs.
Despite its simplicity, DemP consistently achieves superior accuracy-sparsity tradeoffs compared to strong baselines in structured dense-to-sparse training, such as EarlyCrop (DBLP:conf/icml/RachwanZCGAG22) or SNAP (verdenius2020snipit). For instance, when pruning a ResNet-18 on CIFAR-10 or a ResNet-50 on ImageNet, DemP delivers up to a $2.5\%$ improvement in accuracy at $80\%$ sparsity compared to those baselines, while also accelerating training by up to 1.23x on ImageNet. This positions DemP as an efficient method for model compression that balances both accuracy and computational cost. Finally, beyond its strong empirical performance, DemP integrates seamlessly into any training pipeline and can be readily combined with existing pruning techniques.
Our primary contributions are:
@@
    Insights into Neuron Saturation: We explore the mechanisms behind neuron saturation, highlighting the significant roles of stochasticity and key hyperparameters such as learning rate, batch size, and regularization (Section 2 and Appendix B).
@@
    A Simple Structured Pruning Method: We introduce DemP, a dynamic dense-to-sparse training method that promotes neuron death, and prunes them in real-time, by applying decaying regularization and injecting noise early in training (Section 3 and Algorithm 1).
@@
    Training Speedup and High Compression: Extensive experiments on CIFAR-10 and ImageNet show that DemP, despite its simplicity, surpasses strong comparable structured pruning baselines in accuracy-compression tradeoffs, while achieving results comparable to unstructured pruning methods and offering significant training speedup (Section 4 and Appendix G).
1.1 Related Work
Dead Neurons. It is widely recognized that neurons, especially for networks using ReLU activations, can saturate during training (Agarap2018DLRelu; trottier2017parametric; lu2019dying). Continual and reinforcement learning literature linked the accumulation of dead units with plasticity loss (sokar2023dormant; DBLP:conf/icml/LyleZNPPD23; DBLP:journals/corr/AbbasZMWM2023_PlasticityLossInDRL; arXiv2021_DohareRS). Closer to our work, Thesis2018_Evci noted the connection between the dying rate and the learning rate, and derived a pruning technique from it.
relustrikebackMirzadehAMMTSRF highlighted the strategic use of ReLU activations to achieve inference efficiency in transformer-based models and proposed an approach leveraging post-activation sparsity to improve inference speed in large language models.
Structured Pruning. Pruning reduces the size and complexity of neural networks by removing redundant or less important elements, be they neurons or weights (lecun1990optimal). Recent advances such as the Lottery Ticket Hypothesis (LTH) (DBLP:conf/iclr/FrankleC19) demonstrated the existence of subnetworks trainable to performance comparable to their dense counterpart. While unstructured pruning imposes no constraints on which weights to remove, structured methods aim to eliminate entire structures within a network, such as channels, filters, or layers. This results in smaller, faster models that are compatible with existing hardware accelerators and software libraries (wen2016learning; li2016pruning). Structured pruning methods often involve a dense training period followed by sparse fine-tuning (han2015learning). Regularization-based approaches for sparsification after dense training have been widely explored (Louizos2017L0SparseNN; liu2017learning; ye2018rethinking; ding2018auto; Wang2019IncReg; Wang2021GrowReg). The current SOTA for channel pruning at inference, such as Group Fisher (GroupFisherLiuZKZXWCYLZ21), CafeNet-R (CafeNet-RSuYH00Z021), CHIP (ChipSuiYXPZY21) or AutoSlim (AutoSlimYuH2019), all involve either an extended training period or an increased computational budget, negating the potential benefits of pruning for training speedup.
Dynamic Pruning: To address these drawbacks, new methods aim to recover a sparse network under a fixed training budget. Regularization-based methods have been adapted to this setting by leveraging group regularization to influence the training dynamics to yield sparser networks (LessIsMoreZhouAP16; PruneTrainLymCZWSE19), or remarking that batch normalization offers a natural grouping of units under the scale parameter that can be solely regularized (liu2017learning). The LTH inspired multiple methods seeking the winning ticket without iterative pruning. Sparse-to-sparse approaches (SNIPLeeAT19; wang2020grasp)—pruning weights before training—were adapted to work under structured constraints (verdenius2020snipit). Because deep networks suffer increasingly from pruning at initialization (DBLP:conf/icml/FrankleD0C20), dense-to-sparse methods perform pruning at a later stage (EarlyBirdYouL0FWCBWL20; DBLP:conf/icml/RachwanZCGAG22).
A recent trend, Dynamic Sparse Training (DST) (mocanu2018scalable; evci2020rigging), initializes networks sparsely and uses an iteratively updated pruning mask to balance performance and computational efficiency. Initially focusing on unstructured sparsity, DST has recently been adapted for partially-structured sparse networks (Lasby2023srigl; ChaseLFSHWMMPL23), yet still not exceeding 50% node sparsity.
Contextualizing DemP: Within the landscape of pruning techniques, DemP represents an advancement as a dense-to-sparse end-to-end structured method, offering tangible speedup during both training and inference on GPU hardware. It revisits earlier group regularization-based approaches—monitoring activations similar to HuPTT2016network_trimming and leveraging increasing regularization as in (Wang2021GrowReg) on scale parameters (liu2017learning)—by integrating fresh insights into its design to achieve competitiveness with more recent methods while preserving simplicity. Notably, addressing the phenomenon wherein noisier environments can make saturated regions more attractive (see Lemmas  B.2,  B.3 and Appendix D.1), DemP introduces asymmetric noise during optimization to further promote sparsity. Sparsity naturally emerges from the learning dynamics with DemP, simplifying the pruning process and eliminating the need for discrete pruning interventions which could disrupt training. Non-contributing neurons can be pruned as they become inactive, without compromising model performance (as demonstrated in Fig. 7). Recent unstructured pruning approaches manually tailor sparsity distribution across layers (mocanu2018scalable; evci2020rigging), while structured pruning mostly relies on uniform layer-wise pruning (DBLP:conf/icml/RachwanZCGAG22). DemP alleviates this by allowing the learning dynamics to determine sparsity distribution across layers.
DemP eliminates the need for meticulous pruning timing (Wang2021GrowReg; DBLP:conf/icml/RachwanZCGAG22) or heuristic-based pruning schedules (jaxpruner). Acknowledging that early pruning can be detrimental (DBLP:conf/icml/FrankleD0C20; DBLP:conf/icml/RachwanZCGAG22), DemP employs easy-to-tune regularization schedules to influence when units die. Additionally, iterative pruning (verdenius2020snipit) within a single intervention is unnecessary, as units naturally become inactive during early training with DemP.
Choice of Baselines: We highlight and benchmark against strong comparable baselines that use criteria based on gradient flow to evaluate which nodes to prune (verdenius2020snipit; wang2020grasp; DBLP:conf/icml/RachwanZCGAG22). Other methods employing either L0 or L1 regularization on subgroups of parameters to enforce sparsity (liu2017learning; Louizos2017L0SparseNN; You20019GateDecorator; PruneTrainLymCZWSE19) are omitted from the benchmark as they are outperformed by DBLP:conf/icml/RachwanZCGAG22.
@@ 2 Saturating Units: An Analysis
In this section, we offer theoretical insights into the occurrence of dead neurons and explore the impact of different training heuristics and hyperparameters. Given a neural network and a set of $n$ training data samples, we denote by $a^{\ell}_{j}\in\mathbb{R}^{n}$ the vector of activations of the $j$th neuron in layer $\ell$ for each training input. We adopt the following definition of a dead neuron throughout the paper:
Figure 1: Dead neuron accumulation for a ResNet-18 trained on CIFAR-10 with different activation functions and values of the learning rate. We use a negative slope of $\alpha=0.05$ for Leaky ReLU and $\beta=1$ for Swish.
Definition: The $j$-th neuron in layer $\ell$ is inactive if it consistently outputs zero on the entire training set, i.e. $a_{j}^{\ell}=0$. [^1] In practice, especially for non-ReLU activation functions, we will be using the notion of $\epsilon$-inactivity, defined by the condition $|a_{i}^{\ell}|<\epsilon$, for some threshold parameter $\epsilon$. A neuron that becomes and remains inactive during training is considered as dead. [^2] For convolutional layers we treat channels as neurons (see Appendix A).
Many modern architectures use activation functions with a saturation region that typically includes 0 at its boundary. In this case, when a neuron becomes inactive during training, its incoming weights also receive zero—or very small333e.g., due to regularization.—gradients, which makes it difficult for the neuron to recover. In this paper, we mostly work with the Rectified Linear Unit (ReLU) activation function, $\sigma(x)=\max(0,x)$. In this case, the activity of a neuron depends on the sign of the corresponding pre-activation feature.
@@ 2.1 Neurons Die During Training
We begin with some empirical observations. Applying the above criterion (Footnote 1) with threshold parameter $\epsilon=0.01$, we monitor the accumulation of dead neurons during training of a Resnet-18 (DBLP:conf/cvpr/HeZRS16) on CIFAR-10 (krizhevsky2009learning) with the Adam optimizer (DBLP:journals/corr/KingmaB14), with various learning rates and choices of activation functions.
@@
Intuition. Similar to Maxwell’s demon thought experiment (maxwell1872theory), one can picture an asymmetric behavior of how weights can move across the boundary that delimitates the saturated versus the non-saturated state of a unit, akin to the demon’s selective passage of molecules. Neurons, or more specifically their weights, can move freely within the active region, but once they enter the inactive region their movement is impeded, similar to how the demon can trap molecules in a low-energy subspace. In this context, with a random component to their movement, neuron death can be influenced by various factors, e.g., noise from the data, being too close to the border, or taking too large gradient steps. Once the neuron moves into the inactive zone, neurons can only be reactivated if the boundary itself shifts. This asymmetry makes it more likely for neurons to die than to revive.
We formalize this analogy with simple theoretical models in Appendix B. These models are meant to capture the multiplicative (i.e. parameter-dependent) nature of the gradient noise in SGD. Multiplicative noise in stochastic processes is known to cause regions with lower noise magnitude to act as attractors (OksendalSDE_book). Essentially, lower noise regions (such as the inactive region of a saturating unit) tend to retain the system due to reduced impact, while higher noise regions push it away. We explicitly demonstrate this phenomenon in the theoretical framework of Appendix B.
@@ 2.2 Factors Impacting Dying Ratios
Role of regularization. The above intuition suggests that maintaining activations close to zero increases the likelihood of neuron death, aligning them closer to the inactive region (the negative domain for ReLU networks). As illustrated in Fig. 2, training a ResNet-18 on CIFAR-10 (with Adam or SGD+momentum) with a higher L2 or Lasso regularization leads to sparser networks caused by an accumulation of dead units. Section 3 elaborates on the approach with DemP, where we opt to apply regularization exclusively on the scale parameters of the normalization layers, following liu2017learning. This tends to better preserve performance and yields significantly improved accuracy-sparsity tradeoffs, as shown in Fig. 5 in Appendix F.
@@
Figure 2: Left: Increased regularization during training increases the ratio of dead units, as showcased here for a ResNet-18 trained on CIFAR-10. We use $\cdot(\gamma)$ to denote when regularization is applied solely to the scale parameters of the normalization layers. Right: Augmenting training updates with asymmetric Gaussian noise, sampled from $\mathcal{N}(x|0,\sigma^{2})$ and applied to weights of live neurons only, also leads to higher levels of dead unit accumulation, also showcased for a ResNet-18 trained on CIFAR-10 with various values of the Lasso($\gamma$) regularization parameter.
Role of noise. Our simple model suggests a pivotal role played by the noise of the training process in the occurrence of dying neurons. To investigate this in a simple setting, we train small MLPs on a subset of 10,000 images of the MNIST dataset in three different noisy regimes: vanilla SGD, pure SGD noise (obtained by isolating the noisy part of the minibatch gradient), and pure Gaussian noise. As observed in Fig. 9 in Appendix D.1, while pure SGD noise training yields dying ratios comparable to SGD, training with pure Gaussian noise is much less prone to dead neuron accumulation. We hypothesize that this difference is due to the asymmetry of SGD noise: since the gradients are 0 for dead neurons, only (the weights of) live neurons are subject to noisy updates. In contrast, the weights of inactive neurons are updated under Gaussian noise training, which increases their probability of recovery. As we spell out in Section 3, DemP injects Gaussian noise exclusively to weights of live neurons during training, which is effective in encouraging neuron death (see Fig. 2 and Fig. 9, middle plot).
@@
We note that other ways to control the SGD noise variance include varying the learning rate and/or the batch size. Fig. 10 (right) in Appendix D.1 illustrates how increasing the learning rate or reducing the batch size indeed increases dying ratios. However, it proved more challenging to balance accuracy and sparsity by adjusting these hyperparameters, as they also significantly affect overall performance. Instead, DemP maintains these parameters at their default or task-optimized values, facilitating its integration into existing training routines.
@@
Optimizer. We expect the choice of optimizer to affect the final count of dead neurons post-training (e.g., for adaptive optimizers, by altering the effective learning rate per parameter). Notably, we observe a significant discrepancy between using the Adam optimizer (DBLP:journals/corr/KingmaB14) and SGD with momentum (SGDM) (Fig. 2). As also emphasized by DBLP:conf/icml/LyleZNPPD23, we attribute this difference primarily to the specific hyperparameter selection for Adam ($\beta_{1},\beta_{2},\epsilon$), which has a substantial impact on neuron death (further discussed in Appendix E). Results from experiments with both optimizers are presented in the next section.